OpenStudy (anonymous):

Solve \[\sin(x+iy)=0\]

OpenStudy (anonymous):

I got \[y=0, x=2\pi n\]Correct?

hartnn (hartnn):

seems correct.

OpenStudy (anonymous):

But are these the only solutions?

hartnn (hartnn):

they are already infinite....and u know y must be 0 cos sin of imaginary number can't be real

OpenStudy (anonymous):

\[\sin(x+iy)=\sin x \cosh y+i \cos x \sinh y\]

hartnn (hartnn):

*because sin of...

OpenStudy (anonymous):

Cosine can- cosh is a real function. It's imaginary for reals, however

hartnn (hartnn):

cosine can never be imaginary hyperbolic cos can be...i think this is your point...

OpenStudy (anonymous):

Yes. Thanks, I've got it now.